### v2

# From Lab to Field: Real-World Evaluation of an AI-Driven Smart Video Solution to Enhance Community Safety

Abstract
This article adopts and evaluates an AI-enabled Smart Video Solution (SVS) designed to enhance safety in the real world. The system integrates with existing infrastructure camera networks, leveraging recent advancements in AI for easy adoption. Prioritizing privacy and ethical standards, pose based data is used for downstream AI tasks such as anomaly detection. Cloud-based infrastructure and mobile app are deployed, enabling real-time alerts within communities. The SVS employs innovative data representation and visualization techniques, such as the Occupancy Indicator, Statistical Anomaly Detection, Bird’s Eye View, and Heatmaps, to understand pedestrian behaviors and enhance public safety. Evaluation of the SVS demonstrates its capacity to convert complex computer vision outputs into actionable insights for stakeholders, community partners, law enforcement, urban planners, and social scientists. This article presents a comprehensive real-world deployment and evaluation of the SVS, implemented in a community college environment across 16 cameras. The system integrates AI-driven visual processing, supported by statistical analysis, database management, cloud communication, and user notifications. Additionally, the article evaluates the end-to-end latency from the moment an AI algorithm detects anomalous behavior in real-time at the camera level to the time stakeholders receive a notification. The results demonstrate the system’s robustness, effectively managing 16 CCTV cameras with a consistent throughput of 16.5 frames per second (FPS) over a 21-hour period and an average end-to-end latency of 26.76 seconds between anomaly detection and alert issuance.
Introduction
Figure 1: The end-to-end system architecture coupled with notification’s data flow. $S_{1}$ represents the first scene that detects a behavior anomaly. The yellow line shows the notification data flow to the notification service. $S_{n}$ shows the $n^{\text{th}}$ scene where a suspicious object has been detected. The red line shows the notification data flow when detecting object anomalies.
Closed-Circuit Television (CCTV) cameras are widely applied in public spaces, traditionally serving as passive surveillance tools aimed primarily at security and monitoring (Ashby 2017). However, these systems often underutilized the potential of visual data they capture, missing opportunities to contribute to broader civic applications such as urban planning, public health, and resource management (Angelidou et al. 2018). By transforming CCTV networks into Smart Video Solutions (SVS), we can unlock new functionalities that go beyond mere surveillance to enhance community living and environments (Javed et al. 2022).
SVS leverage advancements in AI and machine learning to analyze video data in real-time, providing actionable insights that can improve city infrastructure and community services (Ardabili et al. 2023b). For example, the results can be used to inform urban planners about pedestrian flows for better traffic management (Rashvand et al. 2023), help public health officials monitor social distancing (Usmani et al. 2023), and assist in the efficient allocation of community resources (Dechouniotis et al. 2020). Such applications demonstrate the potential of SVS to transform video data into valuable information for civic improvements.
While traditional research has focused on the development of these systems in controlled environments (Pazho et al. 2023a), it is crucial to transition and evaluate SVS in real-world settings to address challenges such as latency, scalability, and privacy concerns (Ardabili et al. 2023a; Xu et al. 2018). Moving from lab to field ensures that these systems are robust and adaptable, capable of handling the complexities of dynamic urban environments and diverse community needs (Noghre et al. 2023).
The architecture of the SVS is illustrated in Figure 1, which represents the end-to-end design (Ardabili et al. 2023b). It begins with the environment where video data is captured by existing CCTV cameras. The data is then processed by the AI module, where computer vision techniques are applied for human behavior analysis. The processed information is sent to the server module for further data analysis and aggregation (Pazho et al. 2023a). To ensure privacy, metadata is stored in the cloud module, which also hosts a mobile application. This setup allows end-users to receive notifications and access insights on their devices, fostering real-time engagement and decision-making.
This article introduces the deployment and evaluation of an AI-enabled SVS designed to enhance safety and civic functionalities in community spaces such as educational areas, parking lots, and smart city applications. By employing privacy-preserving techniques, the system utilizes pose based data for human behavior analysis, aligning with ethical standards. The SVS applies innovative data visualization tools, including Occupancy Indicators, Statistical Anomaly Detection, Bird’s Eye View, and Heatmaps, to provide insights into pedestrian behaviors and resource allocation, thereby aiding urban planning and enhancing public safety.
The SVS has been implemented in a community college to evaluate its real-world effectiveness. Our findings demonstrate the system’s capability to manage 16 CCTV cameras with a throughput of 16.5 frames per second (FPS) over a 21-hour period, maintaining an average latency of 26.76 seconds for detecting behavioral anomalies and notifying users.
In summary, the contributions of this article are:
  • Adopting an AI-enabled SVS with existing CCTV infrastructure in real world, emphasizing civic applications beyond traditional security measures.
  • Presenting innovative data representation techniques that enhance urban planning and resource allocation by providing insights into pedestrian behaviors.
  • Providing a comprehensive real-world evaluation of the SVS, demonstrating its effectiveness, scalability, and privacy-preserving capabilities in community settings.
Related Works
Recent advancements in AI-driven video systems have leveraged state-of-the-art algorithms and system designs to perform high-level AI tasks using computer vision. Ancilia, for instance, employs a multi-stage computer vision pipeline aimed at enhancing public safety. While we have implemented this system as the baseline for our deployment, it is important to note that the original study reports the results in controlled environment. (Pazho et al. 2023a).
The deployment of SVS in real-world settings is essential to validate their effectiveness and address challenges such as latency, scalability, and privacy. While traditional research has focused on laboratory-based studies for real-time object tracking and anomaly detection (Pazho et al. 2023b; S. et al. 2021), there is a growing recognition of the need to evaluate these systems in practical environments. Real-world evaluation ensures that this technology can adapt to dynamic urban settings and meet diverse community needs.
Recent advancements in machine learning have enabled smart video technology to perform complex data analysis, offering insights for civic applications. For instance, Pramanik et al. demonstrated lightweight models for edge devices to reduce latency in urban environments (Pramanik, Sarkar, and Maiti 2021). Singh et al. introduced an AI-based video technology for fall detection, highlighting the potential for applications beyond traditional safety measures (Singh et al. 2023).
In 2023, the E2E-VSDL method by Gandapur utilized BiGRU and CNN for anomaly detection, achieving high accuracy (Gandapur 2022). Similarly, TAC-Net’s deep learning approach excelled in addressing anomaly scenarios (Huang et al. 2021).
Integrating these systems into smart city applications bridges the gap between theoretical research and practical implementation. Alshammari et al. developed an SVS using surveillance cameras in real-world settings (Alshammari and Rawat 2019), and RV College of Engineering Bengaluru improved object detector accuracy (Franklin and Dabbagol 2020). These efforts highlight the necessity of robust testbed support for evaluating SVS, addressing scalability and privacy challenges (Ma et al. 2019).
The combination of machine learning and IoT devices has transformed urban planning, providing insights for managing congestion and enhancing civic participation. Ardabili et al. introduced an end-to-end system design that integrates Ancilia with an end-user device for direct communication. We have adopted this comprehensive system design for our study’s deployment, allowing us to assess its effectiveness in practical, real-world environments (Ardabili et al. 2023b). Mahdavinejad et al. emphasized machine learning’s role in forecasting urban congestion (Mahdavinejad et al. 2018), while Zanella et al. advocated for accessible urban IoT data (Zanella et al. 2014). These studies illustrate how data-driven approaches contribute to civic applications.
Figure 2: End-to-end detailed system. $C_{0}$ represents the camera ID and for each camera, one AI Module $N_{0}$ including multi-AI-vision-models pipeline is assigned with. All the AI Modules send processed data to one Server Module database, and Server Module will re-identify the human track ID based on the feature extractor data. The statistical analyzer analyzes all the data in the database stored across all the cameras and communicates the results with Cloud Module. Cloud-native services are utilized to host the end users’ applications.
Figure 3: Locations of Cameras on the Campus. A maximum of 16 cameras are used
Despite progress in SVS research, comprehensive testbed support for evaluating real-world system performance remains limited. Our study addresses this gap by deploying a state-of-the-art SVS in a community college, demonstrating its capabilities and potential for enhancing civic applications in public settings.
Software System Features
The AI-based real-time video solution seamlessly integrates with existing CCTV infrastructures, creating a Physical-Cyber-Physical (PCP) system that delivers actionable information to end users. Such system consists of four key components: AI Modules, Server Module, Cloud Module, and End-user Devices, as illustrated in Figure 2. The design of the AI and server modules was inspired by work (Pazho et al. 2023a), while the cloud module and user device components were adapted from another study (Ardabili et al. 2023b).
AI Module/Modules
is a multi-stage pipeline optimized for real-time computer vision tasks. It processes image data in batches of 30 frames, where an object detector identifies objects and flags anomalies (Pazho et al. 2023a). A tracking algorithm then creates tracklets, and pose estimator extracts 2D skeletal data for individuals. Object anomaly detection and behavioral anomaly detection (Noghre, Pazho, and Tabkhi 2024; Pazho et al. 2024) are performed within this module respectively, with alerts sent to end-user devices for real-time response (Pazho et al. 2023b; Yao et al. 2024).
In our feature, YOLOv8 (Jocher, Chaurasia, and Qiu 2023), Bytetrack (Zhang et al. 2022), HRNet (Sun et al. 2019), GEPC (Markovitz et al. 2020a) and OSNet (Zhou et al. 2019) are used for further evaluation. The detailed decision-making process is provided in the supplementary materials for replication and further investigation.
Server Module
centralizes metadata and historical data from AI modules, stored in a MySQL database as outlined in Table 1. It handles global tracking and statistical analysis, using cosine similarity to re-identify individuals across cameras to analyze patterns (Pazho et al. 2023a). The design prioritizes privacy by avoiding raw data transmission to the Cloud Module.
Table 1: Example visualization of the database at Server Module
Record
Time
Camera Class ID
Bounding
Box
Feature Local ID Global ID
Anomaly
Score
00:00:00 1 0 [x, y, w, h] Tensors 15 1001 40
00:00:01 2 0 [x, y, w, h] Tensors 21 1001 40
23:59:59 1 0 [x, y, w, h] Tensors $9$ 1001 40
Cloud Module
leverages cloud-native services for data storage, management, and API generation (Dahunsi, Idogun, and Olawumi 2021). It minimizes lag between anomaly detection and notification, using a rule-based messaging service to send real-time alerts via email, text, or app notifications. A low-latency database supports real-time data access, while APIs generated through an application development kit optimize data retrieval (Mohammed et al. 2022; Ardabili et al. 2023b).
End User Devices
is designed to promptly notify users of detected anomalies via a smartphone application (Ardabili et al. 2023a). The app provides real-time data and analysis, ensuring consistent functionality across different devices and operating systems, thereby enhancing accessibility and delivering information efficiently.
Figure 4: This represents an example of 4 pipelines running in the testbed with IP cameras deployed at different locations.
Deployment and Setup
In lab settings, each advanced software features demonstrates significant potential individually; however, to fully realize its benefits, a comprehensive end-to-end evaluation in a real-world environment is essential. Such an evaluation allows for assessing the system’s load handling and endurance under actual conditions, ensuring that it can perform reliably and effectively when deployed, which might not be evident in a controlled environment.
In our study, an existing network of 16 AXIS IP cameras operating at 30 FPS with 720p resolution across the college campus is applied, covering approximately 35,000 square feet indoors and 60,000 square feet outdoors. Figure 3 visually represents the camera placements. Three cameras monitor outdoor areas, while the remaining 13 oversee various indoor locations, including entry points, vending machines, hallways, and communal spaces. The indoor cameras are mounted at 7 feet 6 inches, and the outdoor cameras at 10 feet 8 inches. Each camera has a varifocal lens with a horizontal field of view of 100 to 36 degrees. The system operates on a dedicated server equipped with a 16-core CPU, 252 GB of memory, and four GPUs, each with 24 GB VRAM.
As shown in Figure 4, a sample output from the Local Node during a scenario with four operational cameras is presented, with privacy-preserving instance segmentation applied.
Applications and Visualizations
This section illustrates how the data collected from real world can be analyzed and visualized to enhance situational awareness in applications such as urban planning, resource allocation, and crowd management.
Descriptive Data
In our descriptive analysis, the Global ID from human feature data serves as the unique identifier for tracking across data streams. We present five key metrics to provide a general overview of traffic flow and occupancy trends: real-time count of people, hourly average number of people per camera per location, total number of people and peak hour analysis over time. These metrics help to monitor occupancy, understand distribution trends, and identify peak traffic times, aiding in resource allocation and emergency planning.
Situational Awareness
Situational awareness is crucial for effective decision-making across various domains. Our study explores four following key visualization techniques:
Occupancy Indicator
provides insight into the number of people in a specific location by comparing current occupancy levels to historical data, which particularly useful in emergencies or public health scenarios. To calculate occupancy, the system compares current detections with historical percentiles, categorizing occupancy as ”Low,” ”Normal,” or ”High.”
Figure 5: Comparing different views of one camera. The graph (a) generated directly from database shows the original camera view. The graph (b) represents the average bird’s eye view coordinates.
(a) Camera 2- Weekday
(b) Camera 2- Weekend
(c) Camera 7- Weekday
(d) Camera 7- Weekend
Figure 6: Heatmap representation of the two cameras during weekday and weekends. Graph (a) and (c) represent the heatmaps during weekdays, while graph (b) and (d) show congestion of the same areas during weekends.
Statistical Anomaly
often indicate unusual traffic flow. By analyzing historical data, the system establishes a baseline, identifying deviations as anomalies. The system calculates such by comparing current detections with historical data, flagging events that exceed two standard deviations from the mean. In our analysis, zero detections were excluded to avoid skewing the data.
Bird’s Eye View
provides an accurate spatial representation, ideal for crowd management and area planning. The system normalizes object dimensions and uses a scale factor to accurately position objects within this view. Figure 5 compares the original and processed Bird’s Eye View from one camera, showing how this technique refines and clarifies spatial data.
Heat map
visually represent data distribution, useful for monitoring congestion and spatial usage. The system generates 2D heatmaps from Bird’s Eye View coordinates, applying Gaussian smoothing for clarity. Figure 6 presents heatmaps for different 24-hour intervals. Graph (a) and (c) show weekday pedestrian movement, while graph (b) and (d) represent the weekend congestions at the same areas. Comparing the density pattern during weekday and weekend, reflect differences in area usage.
The detailed algorithms for all visualization tools are available in the supplementary materials for replication and further investigation.
Community Engagement
To gauge public perception of such solution, we conducted seven rounds of survey studies in six public places across Charlotte, NC, during July and August 2023. A total of 410 respondents participated, providing insights into their concerns about safety and surveillance.
The results indicated that respondents were most concerned about safety in parking lots (54%), public transit (44%), and entertainment venues (42%). Additionally, significant concerns were raised about potential biases and discrimination in current passive surveillance technologies, with these concerns being important or very important to respondents across all demographic groups, highlighting a widespread sensitivity to this issue.  [^1] This study has been previously published, but in adherence to the double-blind review policy of the AAAI conference, the reference will be provided in the camera-ready version upon acceptance.
Interestingly, vulnerable groups, such as females and people of color, were statistically more likely to agree with the statement, ”The presence of video surveillance makes me feel safer.” Furthermore, respondents found our AI-driven video solution more beneficial, with over 50% rating it as very or extremely beneficial, compared to current passive technologies, which 42% rated as not effective or only moderately effective. Our solution also raised fewer privacy concerns, with 66% of respondents reporting no privacy concerns or only moderate concerns. These findings suggest a strong preference for AI-driven surveillance solutions over traditional passive technologies, especially in public places.
System Evaluation and Results
Figure 7: Latency and throughput trends concerning crowd densities during a week-long period for 8 camera nodes
(a) Latency Trends
(b) Throughput Trends
Figure 8: Latency and Throughput trends with respect to crowd densities across different nodes number running in parallelism.
Load Stress Evaluation
This section evaluated the system’s performance under increasing input loads by varying crowd densities from videos. Key metrics such as average latency and throughput were measured as we scaled the system using different numbers of pipeline nodes—one, four, eight, and twelve—to assess scalability and adaptability under increased workloads.
To rigorously assess the system’s performance under varying crowd densities and parallel node operations, we prepared ten videos with density levels ranging from 0 to 9 from CHAD dataset (Pazho et al. 2022). Each video lasted 150 seconds and operated at a frame rate of 30 frames per second. To ensure the statistical validity of our results, we averaged metrics over the central 100 batches in each run, excluding the initial 25 batches for system warm-up and the last 25 batches for cool-down effects.
Figure 8 graphically presents the trends for throughput and latency. The X-axis, labeled ”Density” (count), represents the average number of humans detected per batch (comprising 30 frames) for each input node number in the experiment. With one and four camera nodes, our system maintained latency under 10 seconds, showcasing stability regardless of scene crowdness. However, with eight nodes, latency increased beyond 20 seconds at higher density levels, particularly from level 5 onwards. A notable decrease in latency was observed with 12 nodes, especially at density levels 8 and 9, correlating with processing 108 individuals (12 nodes x 9 density).
Throughput consistently declined linearly with increasing density and node count, reaching a low of 4.56 FPS at the highest density level 9 with 12 nodes. The performance decrease is primarily due to the computational demands of pose estimator (Sun et al. 2019) processing high-density keypoints and is further constrained by CPU computation limits. As node count and density increase, these factors collectively create a bottleneck, leading to the observed decline in throughput and latency.
Real-World Endurance Evaluation
This evaluation assessed the system’s long-term stability using all 16 cameras in the testbed. We conducted extended trials with eight, twelve, and sixteen nodes running continuously for 21 hours and a week-long test using eight nodes to ensure the system’s robustness. This setup aimed to validate the system’s performance in realistic, extended operational conditions.
  • Eight-Node Setup: Covers key indoor areas, including entry points, hallways, and common spaces.
  • Twelve-Node Setup: Adds outdoor monitoring of parking lots.
  • Sixteen-Node Setup: Utilizes all cameras for comprehensive coverage.
As shown in Figure 7, the system’s performance over a week, running with eight cameras, demonstrated consistent latency and throughput results about crowd density. Specifically, the latency ranged from 2.5-7.8s, and throughput varied from 22.3-29.8 FPS. These findings confirm that irrespective of crowd density, the latency increased. At the same time, the FPS decreased, aligning with expectations and underscoring that the length of the videos does not significantly impact the system’s performance.
It is essential to note that the specific data spikes observed within this evaluation could stem from various factors, including network irregularities within the environment (as the input streams originate from IP cameras) or CPU memory usage nearing its limits. However, in light of the comprehensive experiment, these occasional data point spikes remain within acceptable bounds. The system consistently maintained optimal performance levels, showcasing its robustness and capacity to effectively manage and adapt to various operational scenarios.
For configurations with 8, 12, and 16 camera nodes, we observed similar trends over the 21-hour-long experiment with varying performance metrics: average latencies ranged from 2.6-4.8s, 5.3-6.5s, and 6.7-10.5s, respectively, while corresponding throughputs ranged from 28.5-26.5 FPS, 20.5-18 FPS, and 16.5-14.5 FPS. Detailed results with visualizations for 8, 12, and 16 camera nodes are provided in the supplementary materials for replication and further investigation.
(a) Object Anomaly
(b) Behavior Anomaly
Figure 9: Represent latency rug plot with different anomaly scenarios. 50 data points for each camera number
Physical-Cyber-Physical Evaluation (Anomaly detection)
This section evaluates the system’s Physical-Cyber-Physical (PCP) latency, representing the end-to-end time from when an anomaly appears to when the end user receives a notification. We assessed both object and behavioral anomaly detection (Markovitz et al. 2020b) using four, eight, twelve, and sixteen AI pipeline nodes reading input frames from the same camera node, conducting twenty experiments per anomaly type and collecting 400 data samples. The latency measures the time from detecting anomalies, such as high-priority objects or dangerous behaviors like fighting, to alerting the end user. A three-second transmission delay from cameras to AI Nodes was observed, slightly higher than other IP systems, due to our unique network setup. This delay is included in the overall PCP measurements, though manual recording of latency data may introduce minor discrepancies.
The results indicate that increasing the number of cameras significantly impacts PCP latency. For example, object anomaly latency increased from 4.7 seconds with four cameras to 19.68 seconds with sixteen, while behavioral anomaly latency rose from 9.34 seconds to 26.76 seconds. This increase is attributed to system traffic, multitasking overhead, and the processing capacity of the system.
Figure 9 displays the distribution of 400 data points from our PCP experiments. Graph (a) shows the distribution for object anomaly detection, with less variation in the four-camera setup compared to the sixteen-camera setup. Graph (b) illustrates the distribution for behavioral anomaly detection, where the data is more scattered with an increased number of cameras. The standard deviation for object detection latency increased from 0.65 to 1.12 seconds as the camera count increased, indicating a slight broadening of latency due to system complexity and resource contention.
This evaluation demonstrates the system’s capability to enhance safety and operational efficiency in real-world scenarios. The ability to notify end users within approximately 20 seconds is crucial for managing anomalies in public spaces, aiding in the prevention of potential dangers and improving resource allocation.
Overall, these real-world evaluations allows research how these key parameters, such as the number of cameras and crowd density, could affect system performance, such as latency and throughput in different public safety scenarios in case of emergency, is crucial in decision making and taking the proper actions. For example, retail and transport hubs could analyze foot traffic and crowd density, offering insights into peak hours and emergency response planning (Hosseini et al. 2024).
Conclusion
This paper presented a comprehensive evaluation of an AI-based real-time video solution, integrated with existing CCTV infrastructures, to enhance situational awareness through object and behavioral anomaly detection. The system was tested across various configurations, including up to sixteen camera nodes, to assess its performance under real-world conditions.
Our evaluation highlighted the system’s ability to maintain operational efficiency with increasing node counts, despite a corresponding rise in PCP latency. Object and behavioral anomaly detection latencies ranged from 4.7 seconds with four cameras to 26.76 seconds with sixteen, reflecting the system’s scalability and robustness. Through extensive testing, the system demonstrated its capability to deliver timely notifications to end users, crucial for managing anomalies in public spaces. This work underscores the importance of real-world evaluations in optimizing AI-driven video solutions, ensuring they meet the demands of dynamic environments and contribute to enhanced public safety and operational efficiency.

### v3

@@
 The system integrates with existing infrastructure camera networks, leveraging recent advancements in AI for easy adoption.
-Prioritizing privacy and ethical standards, pose based data is used for downstream AI tasks such as anomaly detection.
-Cloud-based infrastructure and mobile app are deployed, enabling real-time alerts within communities.
+Prioritizing privacy and ethical standards, pose-based data is used for downstream AI tasks such as anomaly detection.
+A Cloud-based infrastructure and a mobile app are deployed, enabling real-time alerts within communities.
 The SVS employs innovative data representation and visualization techniques, such as the Occupancy Indicator, Statistical Anomaly Detection, Bird’s Eye View, and Heatmaps, to understand pedestrian behaviors and enhance public safety.
 Evaluation of the SVS demonstrates its capacity to convert complex computer vision outputs into actionable insights for stakeholders, community partners, law enforcement, urban planners, and social scientists.
-This article presents a comprehensive real-world deployment and evaluation of the SVS, implemented in a community college environment across 16 cameras.
+This article presents a comprehensive real-world deployment and evaluation of the SVS, implemented in a community college environment with 16 cameras.
 The system integrates AI-driven visual processing, supported by statistical analysis, database management, cloud communication, and user notifications.
@@
 The results demonstrate the system’s robustness, effectively managing 16 CCTV cameras with a consistent throughput of 16.5 frames per second (FPS) over a 21-hour period and an average end-to-end latency of 26.76 seconds between anomaly detection and alert issuance.
-Introduction
-Figure 1: The end-to-end system architecture coupled with notification’s data flow.
+keywords:
+Smart Video Solution, Computer Vision, Case Study, Public Safety, Anomaly Detection
+journal: Internet of Things
+[…]
+[label1]organization=Department of Electrical and Computer Engineering, University of North Carolina at Charlotte,city=Charlotte, state=NC, country=USA
+[…]
+[label2]organization=Department of Public Policy, University of North Carolina at Charlotte,city=Charlotte, state=NC, country=USA
+1 Introduction
+Closed-Circuit Television (CCTV) cameras are widely applied in public spaces, traditionally serving as passive surveillance tools aimed primarily at security and monitoring [1].
+However, these systems often underutilizing the potential of visual data they capture, missing opportunities to contribute to broader civic applications such as urban planning, public health, and resource management [2].
+By transforming CCTV networks into Smart Video Solutions (SVS), we can unlock new functionalities that go beyond mere surveillance to enhance community living and environments [3].
+SVS utilizes advanced AI and machine learning techniques to conduct real-time analysis of video data, thereby generating actionable insights that enhance urban infrastructure and community services [4].
+Specifically, these insights can inform urban planners about pedestrian flows to improve traffic management [5], enable public health officials to monitor social distancing [6], and facilitate the efficient allocation of community resources [7].
+Collectively, these applications underscore the significant potential of SVS to transform video data into valuable information for civic improvements.
+While previous research has focused on developing SVS in controlled environments [8], transitioning and evaluating these systems in real-world settings is crucial to address challenges such as latency, scalability, and privacy concerns [9, 10].
+This shift from controlled laboratory settings to diverse real-world environments ensures that SVS is robust against real-world variability and adaptable to changing conditions, enabling it to effectively handle the complexities of dynamic urban environments and diverse community needs [11].
+This study primarily focuses on evaluating of such a system in real-world settings, leveraging both the AI [8] and application [4] components from the earlier State-Of-The-Art (SOTA) research that emphasized the development of such a system.
+Through this real-world evaluation, we are assessing the system’s impact in community settings and its broader civic applications.
+Figure 1: The end-to-end system architecture is coupled with the notification’s data flow.
 $S_{1}$ represents the first scene that detects a behavior anomaly.
@@
 The red line shows the notification data flow when detecting object anomalies.
-Closed-Circuit Television (CCTV) cameras are widely applied in public spaces, traditionally serving as passive surveillance tools aimed primarily at security and monitoring (Ashby 2017).
-However, these systems often underutilized the potential of visual data they capture, missing opportunities to contribute to broader civic applications such as urban planning, public health, and resource management (Angelidou et al.
-2018).
-By transforming CCTV networks into Smart Video Solutions (SVS), we can unlock new functionalities that go beyond mere surveillance to enhance community living and environments (Javed et al.
-2022).
-SVS leverage advancements in AI and machine learning to analyze video data in real-time, providing actionable insights that can improve city infrastructure and community services (Ardabili et al.
-2023b).
-For example, the results can be used to inform urban planners about pedestrian flows for better traffic management (Rashvand et al.
-2023), help public health officials monitor social distancing (Usmani et al.
-2023), and assist in the efficient allocation of community resources (Dechouniotis et al.
-2020).
-Such applications demonstrate the potential of SVS to transform video data into valuable information for civic improvements.
-While traditional research has focused on the development of these systems in controlled environments (Pazho et al.
-2023a), it is crucial to transition and evaluate SVS in real-world settings to address challenges such as latency, scalability, and privacy concerns (Ardabili et al.
-2023a; Xu et al.
-2018).
-Moving from lab to field ensures that these systems are robust and adaptable, capable of handling the complexities of dynamic urban environments and diverse community needs (Noghre et al.
-2023).
-The architecture of the SVS is illustrated in Figure 1, which represents the end-to-end design (Ardabili et al.
-2023b).
+The architecture of the SVS is illustrated in Fig.
+1, which represents the end-to-end design [4].
 It begins with the environment where video data is captured by existing CCTV cameras.
 The data is then processed by the AI module, where computer vision techniques are applied for human behavior analysis.
-The processed information is sent to the server module for further data analysis and aggregation (Pazho et al.
-2023a).
+The processed information is sent to the server module for further data analysis and aggregation [8].
 To ensure privacy, metadata is stored in the cloud module, which also hosts a mobile application.
 This setup allows end-users to receive notifications and access insights on their devices, fostering real-time engagement and decision-making.
+Previous studies have mainly focused on the development of SVS either from the technical or the application perspectives.
+However, their approaches were mainly limited to lab settings, which lacked real-world validations.
+Deploying these systems in the real world, introduce new challenges such as network latency, integration with the existing infrastructure, and scalability.
+Moreover, the main scope of SVS focuses on security applications like detecting behavior anomalies or detecting fire hazards.
+Generally, there is less debate on privacy in these applications, where the focus is on immediate identification of security threats.
+However, when there is a need for continuous monitoring of public spaces in broader civic applications, privacy concerns become significantly more important.
+This further underscores the need to assess real-world performance to ensure that the system operates effectively while complying with ethical aspects.
+In this study, we bridge the gap in deploying and evaluating an SVS in the real-world setup and go beyond security applications, to show the scalability and robustness in broader civic applications.
 This article introduces the deployment and evaluation of an AI-enabled SVS designed to enhance safety and civic functionalities in community spaces such as educational areas, parking lots, and smart city applications.
-By employing privacy-preserving techniques, the system utilizes pose based data for human behavior analysis, aligning with ethical standards.
+By employing privacy-preserving techniques, the system utilizes pose-based data for human behavior analysis, aligning with ethical standards.
 The SVS applies innovative data visualization tools, including Occupancy Indicators, Statistical Anomaly Detection, Bird’s Eye View, and Heatmaps, to provide insights into pedestrian behaviors and resource allocation, thereby aiding urban planning and enhancing public safety.
@@
 In summary, the contributions of this article are:
-  • Adopting an AI-enabled SVS with existing CCTV infrastructure in real world, emphasizing civic applications beyond traditional security measures.
+  • 1.
+    Adopting an AI-enabled SVS with existing CCTV infrastructure in the real world, emphasizing civic applications beyond traditional security measures.
+  • 2.
     Presenting innovative data representation techniques that enhance urban planning and resource allocation by providing insights into pedestrian behaviors.
+  • 3.
     Providing a comprehensive real-world evaluation of the SVS, demonstrating its effectiveness, scalability, and privacy-preserving capabilities in community settings.
-Related Works
-Recent advancements in AI-driven video systems have leveraged state-of-the-art algorithms and system designs to perform high-level AI tasks using computer vision.
-Ancilia, for instance, employs a multi-stage computer vision pipeline aimed at enhancing public safety.
-While we have implemented this system as the baseline for our deployment, it is important to note that the original study reports the results in controlled environment.
-(Pazho et al.
-2023a).
-The deployment of SVS in real-world settings is essential to validate their effectiveness and address challenges such as latency, scalability, and privacy.
-While traditional research has focused on laboratory-based studies for real-time object tracking and anomaly detection (Pazho et al.
-2023b; S.
-et al.
-2021), there is a growing recognition of the need to evaluate these systems in practical environments.
-Real-world evaluation ensures that this technology can adapt to dynamic urban settings and meet diverse community needs.
+The remainder of this paper is structured as follows.
+Section 2 provides a comprehensive review of relevant literature in the field.
+Section 3 outlines the key features and architecture of our system.
+In Section 4, we detail the implementation methodology and evaluation framework.
+Section 5 presents our data visualization tools and techniques, while Section 6 examines community engagement and its impact.
+Section 7 describes our experimental setup and discusses the results.
+Finally, Section 8 summarizes our findings and discusses future research directions.
+2 Related Works
+Recent advancements in AI-driven video systems have harnessed state-of-the-art algorithms and system designs to execute high-level AI tasks through computer vision.
+For example, Ancilia [8] utilizes a multi-stage computer vision pipeline specifically engineered to enhance public safety.
+Although we adopted this system as the baseline for our deployment, the original study reports results exclusively from a controlled environment.
+This limitation underscores the necessity of real-world evaluations to confirm the system’s practical efficacy.
+Deploying SVS systems in real-world environments is critical to validate their performance and tackle key challenges, including latency, scalability, and privacy concerns.
+Historically, research has concentrated on laboratory-based studies for real-time object tracking and anomaly detection [12, 13].
+However, there is increasing acknowledgment of the need to assess these systems in dynamic, real-world settings.
+Such evaluations are vital to ensure SVS technologies can adapt to the complexities of urban landscapes and fulfill diverse community requirements.
 Recent advancements in machine learning have enabled smart video technology to perform complex data analysis, offering insights for civic applications.
 For instance, Pramanik et al.
-demonstrated lightweight models for edge devices to reduce latency in urban environments (Pramanik, Sarkar, and Maiti 2021).
+demonstrated lightweight models for edge devices to reduce latency in urban environments [14].
 Singh et al.
-introduced an AI-based video technology for fall detection, highlighting the potential for applications beyond traditional safety measures (Singh et al.
-2023).
-In 2023, the E2E-VSDL method by Gandapur utilized BiGRU and CNN for anomaly detection, achieving high accuracy (Gandapur 2022).
-Similarly, TAC-Net’s deep learning approach excelled in addressing anomaly scenarios (Huang et al.
-2021).
+introduced an AI-based video technology for fall detection, highlighting the potential for applications beyond traditional safety measures [15].
+In 2023, the E2E-VSDL method by Gandapur utilized BiGRU and CNN for anomaly detection, achieving high accuracy [16].
+Similarly, TAC-Net’s deep learning approach excelled in addressing anomaly scenarios [17].
 Integrating these systems into smart city applications bridges the gap between theoretical research and practical implementation.
 Alshammari et al.
-developed an SVS using surveillance cameras in real-world settings (Alshammari and Rawat 2019), and RV College of Engineering Bengaluru improved object detector accuracy (Franklin and Dabbagol 2020).
-These efforts highlight the necessity of robust testbed support for evaluating SVS, addressing scalability and privacy challenges (Ma et al.
-2019).
+developed an SVS using surveillance cameras in real-world settings [18], and RV College of Engineering Bengaluru improved object detector accuracy [19].
+These efforts highlight the necessity of robust testbed support for evaluating SVS, addressing scalability and privacy challenges [20].
+Y.
+Yuan et al.
+[21] introduced an innovative decentralized framework for Cyber-Physical systems, demonstrating its effectiveness in optimizing traffic flow and reducing pedestrian waiting times in real-world scenarios.
+While [22] presents an impressive self-aware Cyber-Physical system, it notably lacks comprehensive data on system latency—a crucial metric in testbed environments.
+Recent research has increasingly focused on optimizing large-scale Smart Vehicle systems, with significant attention given to enhancing data transformation, communication efficiency, and the integration of blockchain technology across various domains [23].
+These developments demonstrate a shift beyond traditional model performance metrics to address comprehensive system scalability and real-world implementation challenges.
 The combination of machine learning and IoT devices has transformed urban planning, providing insights for managing congestion and enhancing civic participation.
@@
 introduced an end-to-end system design that integrates Ancilia with an end-user device for direct communication.
-We have adopted this comprehensive system design for our study’s deployment, allowing us to assess its effectiveness in practical, real-world environments (Ardabili et al.
-2023b).
+We have adopted this comprehensive system design for our study’s deployment, allowing us to assess its effectiveness in practical, real-world environments [4].
 Mahdavinejad et al.
-emphasized machine learning’s role in forecasting urban congestion (Mahdavinejad et al.
-2018), while Zanella et al.
-advocated for accessible urban IoT data (Zanella et al.
-2014).
+emphasized machine learning’s role in forecasting urban congestion [24], while Zanella et al.
+advocated for accessible urban IoT data [25].
 These studies illustrate how data-driven approaches contribute to civic applications.
+Despite significant advances in Smart Vehicle Systems research, there remains a critical gap in the comprehensive testbed evaluation of real-world SVS performance.
+To address this limitation, we have developed and implemented an SVS system within a public community college setting, providing a practical demonstration of these technologies in a real-world environment.
+The following sections detail our implementation, findings, and their implications for future SVS deployments.
+3 Software System Features
 Figure 2: End-to-end detailed system.
-$C_{0}$ represents the camera ID and for each camera, one AI Module $N_{0}$ including multi-AI-vision-models pipeline is assigned with.
-All the AI Modules send processed data to one Server Module database, and Server Module will re-identify the human track ID based on the feature extractor data.
-The statistical analyzer analyzes all the data in the database stored across all the cameras and communicates the results with Cloud Module.
-Cloud-native services are utilized to host the end users’ applications.
-Figure 3: Locations of Cameras on the Campus. A maximum of 16 cameras are used
-Despite progress in SVS research, comprehensive testbed support for evaluating real-world system performance remains limited.
-Our study addresses this gap by deploying a state-of-the-art SVS in a community college, demonstrating its capabilities and potential for enhancing civic applications in public settings.
-Software System Features
+$C_{0}$ represents the camera ID and for each camera, one AI Module $N_{0}$ including a multi-AI-vision-models pipeline is assigned.
+All the AI Modules send processed data to one Server Module database, and the Server Module will re-identify the human track ID based on the feature extractor data.
+The statistical analyzer analyzes all the data stored in the database across all the cameras
+and communicates the results with the Cloud Module. Cloud-native services are utilized to host the end users’ applications.
+Algorithm 1 SVS Algorithm
+1:Initialization:
+2:  Initialize AI modules, Server Module, and Cloud Module.
+3:for each camera $c_{i}$ in the set of cameras do
+4:  while stream from $c_{i}$ is active do
+5:   Step 1: Object Detection
+6:    Fetch a batch of 30 frames of rgbrgb data from $c_{i}$ as initial input
+7:    Apply the object detector to identify humans and objects of interest.
+8:    Record bounding boxes $B$ with associated confidence scores.
+9:   if objects of interest confidence score in $B\geq\textit{object\_threshold}$ then
+10:    Trigger an alert for the detected object of interest to end-user device(s).
+11:   else
+12:    Continue
+13:   end if
+14:   Step 2: Tracking
+15:    Input $B$ to the tracker to assign unique IDs and generate tracklets $T$.
+16:   Step 3: Pose Estimation
+17:    For each tracklet in $T$, extract the corresponding input rgbrgb data.
+18:    Apply the pose estimator to compute skeletal keypoints $K$.
+19:   Step 4: Anomaly Score Prediction
+20:    Organize input keypoints into movement sequences.
+21:    Compute anomaly scores $\alpha$ for the batch of keypoints.
+22:   Step 5: Feature Extraction & Cross-Camera Identification
+23:    Extract feature $F$ from input rgbrgb for cross-camera identification.
+24:    Transmit $F$ and $\alpha$ to the Server Module for global analyze.
+25:   Step 6: Behavior Anomaly Detection
+26:    Transmit $\alpha$ to the Server Module for behavior anomaly detection.
+27:   if $\alpha\geq\textit{threshold}$ for three consecutive batches then
+28:    Trigger an alert to the end-user device(s).
+29:   else
+30:    Continue
+31:   end if
+32:   Step 7: Data Processing at Server Module
+33:    Process $F$, anomaly scores $\alpha$, and related data for statistical analysis.
+34:   Step 8: Data Transmission
+35:    Transmit analyzed results to the Cloud and output results to end-user device(s) with different visualizations.
+36:  end while
+37:end for
 The AI-based real-time video solution seamlessly integrates with existing CCTV infrastructures, creating a Physical-Cyber-Physical (PCP) system that delivers actionable information to end users.
-Such system consists of four key components: AI Modules, Server Module, Cloud Module, and End-user Devices, as illustrated in Figure 2.
-The design of the AI and server modules was inspired by work (Pazho et al.
-2023a), while the cloud module and user device components were adapted from another study (Ardabili et al.
-2023b).
-AI Module/Modules
-is a multi-stage pipeline optimized for real-time computer vision tasks.
-It processes image data in batches of 30 frames, where an object detector identifies objects and flags anomalies (Pazho et al.
-2023a).
-A tracking algorithm then creates tracklets, and pose estimator extracts 2D skeletal data for individuals.
-Object anomaly detection and behavioral anomaly detection (Noghre, Pazho, and Tabkhi 2024; Pazho et al.
-2024) are performed within this module respectively, with alerts sent to end-user devices for real-time response (Pazho et al.
-2023b; Yao et al.
-2024).
-In our feature, YOLOv8 (Jocher, Chaurasia, and Qiu 2023), Bytetrack (Zhang et al.
-2022), HRNet (Sun et al.
-2019), GEPC (Markovitz et al.
-2020a) and OSNet (Zhou et al.
-2019) are used for further evaluation.
-The detailed decision-making process is provided in the supplementary materials for replication and further investigation.
-Server Module
-centralizes metadata and historical data from AI modules, stored in a MySQL database as outlined in Table 1.
-It handles global tracking and statistical analysis, using cosine similarity to re-identify individuals across cameras to analyze patterns (Pazho et al.
-2023a).
-The design prioritizes privacy by avoiding raw data transmission to the Cloud Module.
+Such a system consists of four key components: AI Modules, Server Module, Cloud Module, and End-user Devices, as illustrated in Fig.
+2.
+Multiple camera streams can be fed into separate AI Modules in parallel.
+The object detector identifies bounding boxes for humans and other objects of interest, then forwards this bounding-box data to a tracker, which assigns unique IDs.
+Next, the tracked bounding boxes, including associated pixel data, are passed to a pose estimator to determine key skeletal joints.
+These skeletal data joints, organized into movement sequences, are subsequently processed by an anomaly detector to produce an anomaly score for the scene.
+All previously processed data are then sent to a feature-extraction stage to generate unique signatures for cross-camera identification, a task carried out by the server module.
+Finally, additional statistical data are analyzed and transmitted to the cloud module for end-user applications.
+The design of the AI and server modules was inspired by prior work [8], while the cloud module and user device components were adapted from another study [4].
+Algorithm 1 provides more details regarding the complete SVS.
+3.1 AI Module/Modules
+AI Module/Modules is a multi-stage pipeline optimized for real-time computer vision tasks.
+It processes image data in batches of 30 frames, in which the object detector identifies objects and the anomaly detector flags the behavior [8].
+Object anomaly detection and behavioral anomaly detection [26, 27] are performed within this module respectively, with alerts sent to end-user devices for real-time response [12, 28].
+One fundamental experiment was conducted to identify the optimal algorithm configuration for the AI node [8], specifically tailored to address real-world security requirements before going into further research.
+As highlighted by Pazho et al.
+[8], the GEPC method [29] is particularly effective for behavioral anomaly detection.
+ByteTrack [30] and OSNet [31] excel in pedestrian tracking and human re-identification, while the recent YOLOv8 [32] demonstrates superior performance in object detection and enhances human pose estimation through a top-down approach.
+Such evaluation utilized the DukeMTMC dataset [33], with one-minute video clips of varying crowd densities.
+Initial validation was conducted on a laboratory server using default weights and parameters from original repositories, measuring throughput, latency, detection outputs, and computational resource utilization.
+The testing protocol employed two high-density scenarios, labeled ”Extreme” and ”Heavy,” each comprising 60-second videos at 60 frames per second, yielding 120 inference batches.
+Statistical validity was ensured by analyzing the middle 80 batches exclusively, eliminating potential bias from warm-up and cool-down periods in the performance metrics.
+Four distinct methods underwent testing:
+  • 1.
+    YOLOv5+HRNet: The original system method
+  • 2.
+    YOLOv8pose: A top-down human pose estimation method.
+  • 3.
+    YOLOv8pose-p6: The most complex model of YOLOv8 for pose estimation.
+  • 4.
+    YOLOv8+HRNet: Object detection method with human pose estimation as a bottom-up method.
+Table 1: Performance comparison among different system configurations. Data collected using lab server with a single node.
+Crowd
+Density
+Method
+Latency
+FPS
+Detections
+GPU
+Memory
+CPU
+Memory
+Total
+FLOPs
+Extreme ($\sim$50 detects per second) YOLOv5+HRNet 14 14.82 1033 5.3 19.1 126.08
+YOLOv8pose-p6 3.7 26.39 911 12.2 12.68 1067.38
+YOLOv8pose 2.2 49.07 679 6.9 12.38 264.18
+YOLOv8+HRNet 9.09 19 1217 5.5 19.3 182.18
+Heavy ($\sim$20 detects per second) YOLOv5+HRNet 3.47 81.2 237 5.0 15.28 126.08
+YOLOv8pose-p6 3.54 26.45 358 11.9 12.4 1067.38
+YOLOv8pose 2.6 45.11 308 6.9 12.3 264.18
+YOLOv8+HRNet 2.76 67.63 342 5.5 15.4 182.18
+Table 1 evaluates system performance across varying crowd densities using four key metrics: average latency and throughput, total detection count, average GPU and CPU memory utilization, and the AI models’ total FLOPs.
+The combined YOLOv8+HRNet architecture demonstrated superior detection capabilities, particularly in challenging ”Extreme” and ”Heavy” crowd density scenarios, validating its robustness for real-world applications.
+For a comprehensive evaluation, our pipeline integrates YOLOv8 [32] for object detection, ByteTrack [30] for tracking, HRNet [34] for pose estimation, GEPC [35] for behavioral analysis, and OSNet [31] for re-identification.
+Table 2: Accuray of different model weights implemented in the system
+Model Dataset Metric Value
+YoloV8l [32] COCO mAP 52.9
+ByteTrack [30] MOT17 MOTA 80.3
+HRNet_w32 [34] COCO AP 0.744
+GEPC [29] NTU60 AUC-ROC 0.85
+OSNet [31] OSNet R1 94.8
+Importantly, all accuracy metrics in subsequent experiments are derived directly from the pre-trained model weights as published in each respective repository, without any additional fine-tuning or optimization.
+This approach ensures reproducibility and maintains consistency with the established benchmarks reported in the original publications.
+Table 2 summarizes the accuracy metrics of each model that is used for experiments on the existing well-established datasets.
+3.2 Server Module
 Table 3: Example visualization of the database at Server Module
-Record
-Time
-Camera Class ID
-Bounding
-Box
-Feature Local ID Global ID
-Anomaly
-Score
+Record Time Camera Class ID Bounding Box Feature Local ID Global ID Anomaly Score
 00:00:00 1 0 [x, y, w, h] Tensors 15 1001 40
 00:00:01 2 0 [x, y, w, h] Tensors 21 1001 40
-23:59:59 1 0 [x, y, w, h] Tensors $9$ 1001 40
-Cloud Module
-leverages cloud-native services for data storage, management, and API generation (Dahunsi, Idogun, and Olawumi 2021).
+23:59:59 1 0 [x, y, w, h] Tensors 9999 1001 40
+Server Module centralizes metadata and historical data from AI modules, stored in a MySQL database as outlined in Table 3.
+It handles global tracking and statistical analysis, using cosine similarity to re-identify individuals across cameras and analyze patterns [8].
+The design prioritizes privacy by avoiding raw data transmission to the Cloud Module.
+3.3 Cloud Module
+Cloud Module leverages cloud-native services for data storage, management, and API generation [36].
 It minimizes lag between anomaly detection and notification, using a rule-based messaging service to send real-time alerts via email, text, or app notifications.
-A low-latency database supports real-time data access, while APIs generated through an application development kit optimize data retrieval (Mohammed et al.
-2022; Ardabili et al.
-2023b).
-End User Devices
-is designed to promptly notify users of detected anomalies via a smartphone application (Ardabili et al.
-2023a).
+A low-latency database supports real-time data access, while APIs generated through an application development kit optimize data retrieval [37, 4].
+3.4 End User Devices
+End User Devices are designed to promptly notify users of detected anomalies via a smartphone application [9].
 The app provides real-time data and analysis, ensuring consistent functionality across different devices and operating systems, thereby enhancing accessibility and delivering information efficiently.
-Figure 4: This represents an example of 4 pipelines running in the testbed with IP cameras deployed at different locations.
-Deployment and Setup
-In lab settings, each advanced software features demonstrates significant potential individually; however, to fully realize its benefits, a comprehensive end-to-end evaluation in a real-world environment is essential.
+4 Deployment and Setup
+Figure 3: Locations of Cameras on the Campus. A maximum of 16 cameras are used
+In lab settings, each advanced software feature demonstrates significant potential individually; however, to fully realize its benefits, a comprehensive end-to-end evaluation in a real-world environment is essential.
 Such an evaluation allows for assessing the system’s load handling and endurance under actual conditions, ensuring that it can perform reliably and effectively when deployed, which might not be evident in a controlled environment.
-In our study, an existing network of 16 AXIS IP cameras operating at 30 FPS with 720p resolution across the college campus is applied, covering approximately 35,000 square feet indoors and 60,000 square feet outdoors.
-Figure 3 visually represents the camera placements.
-Three cameras monitor outdoor areas, while the remaining 13 oversee various indoor locations, including entry points, vending machines, hallways, and communal spaces.
-The indoor cameras are mounted at 7 feet 6 inches, and the outdoor cameras at 10 feet 8 inches.
-Each camera has a varifocal lens with a horizontal field of view of 100 to 36 degrees.
-The system operates on a dedicated server equipped with a 16-core CPU, 252 GB of memory, and four GPUs, each with 24 GB VRAM.
-As shown in Figure 4, a sample output from the Local Node during a scenario with four operational cameras is presented, with privacy-preserving instance segmentation applied.
-Applications and Visualizations
-This section illustrates how the data collected from real world can be analyzed and visualized to enhance situational awareness in applications such as urban planning, resource allocation, and crowd management.
-Descriptive Data
+Our study utilizes an existing surveillance network of 16 AXIS IP cameras, operating at 30 FPS with 720p resolution, across the campus, monitoring approximately 95,000 square feet of combined indoor and outdoor space.
+As shown in Fig.
+3, thirteen indoor cameras are mounted and angled slightly downward to capture hallways, entryways, vending areas, and other common spaces, while three outdoor cameras, installed at 10 ft 8 in, oversee building exteriors, walkways, and surrounding grounds.
+All cameras feature varifocal lenses offering a 36–100° horizontal field of view, allowing flexible adjustments to accommodate wide-area observation.
+The system runs on a dedicated server with a 16-core CPU, 252 GB RAM, and four 24 GB VRAM GPUs.
+The software components, encompassing the AI module and the Server Module as introduced in Section 3, are integrated into a dedicated server infrastructure.
+This server is interconnected within the local network of the testbed to enhance security measures.
+Video feeds from all cameras are transmitted utilizing RTSP.
+Data processed by each AI module is subsequently relayed to the database housed within the Server Module via TCP.
+The Cloud module server periodically aggregates data from the Server Module’s database for advanced data management purposes.
+The resultant outputs are then disseminated and updated across end-user applications
+As shown in Fig. 4, a sample visualized output from the AI module during a scenario with four operational cameras is presented, with privacy-preserving instance segmentation applied.
+Figure 4: This represents an example of four pipelines running in the testbed, with IP cameras deployed at different locations.
+5 Applications and Visualizations
+This section illustrates how the data collected from the real world can be analyzed and visualized to enhance situational awareness in different civic applications.
+For example, visualizing pedestrian flow can inform the design of public spaces like intersection turnarounds to reduce congestion, while tracking real-time movement patterns helps prevent overcrowding and ensures safety in public events.
+These applications showcase how data-driven insights can lead to more efficient, safe, and well-planned communities.
+5.1 Descriptive Data
 In our descriptive analysis, the Global ID from human feature data serves as the unique identifier for tracking across data streams.
-We present five key metrics to provide a general overview of traffic flow and occupancy trends: real-time count of people, hourly average number of people per camera per location, total number of people and peak hour analysis over time.
-These metrics help to monitor occupancy, understand distribution trends, and identify peak traffic times, aiding in resource allocation and emergency planning.
-Situational Awareness
-Situational awareness is crucial for effective decision-making across various domains. Our study explores four following key visualization techniques:
-Occupancy Indicator
-provides insight into the number of people in a specific location by comparing current occupancy levels to historical data, which particularly useful in emergencies or public health scenarios.
+We present five key metrics to provide a general overview of traffic flow and occupancy trends: real-time count of people, hourly average number of people per camera per location, total number of people, and peak hour analysis over time.
+These metrics help monitor occupancy, understand distribution trends, and identify peak traffic times, aiding in resource allocation and emergency planning.
+5.2 Situational Awareness
+Situational awareness is crucial for effective decision-making across various domains. Our study explores the four following key visualization techniques:
+5.2.1 Occupancy Indicator
+The Occupancy Indicator provides insight into the number of people in a specific location by comparing current occupancy levels to historical data, which is particularly useful in emergencies or public health scenarios.
 To calculate occupancy, the system compares current detections with historical percentiles, categorizing occupancy as ”Low,” ”Normal,” or ”High.”
-Figure 5: Comparing different views of one camera. The graph (a) generated directly from database shows the original camera view. The graph (b) represents the average bird’s eye view coordinates.
-(a) Camera 2- Weekday
-(b) Camera 2- Weekend
-(c) Camera 7- Weekday
-(d) Camera 7- Weekend
-Figure 6: Heatmap representation of the two cameras during weekday and weekends.
-Graph (a) and (c) represent the heatmaps during weekdays, while graph (b) and (d) show congestion of the same areas during weekends.
-Statistical Anomaly
-often indicate unusual traffic flow.
+Figure 5: Comparing the Occupancy Indicator of the same camera during the same hour on two different days
+. Graph (1) shows the Weekend. Graph (2) represents weekday.
+Algorithm 2 describes a method for calculating and categorizing real-time occupancy levels across multiple camera locations.
+Operating on a continuously updated data frame (df), the algorithm processes each camera location sequentially.
+For each camera ID, it performs occupancy analysis at 5-second intervals throughout the specified record time.
+During each interval, the algorithm computes the current occupancy by determining the number of unique global IDs present in the data frame, effectively tracking the distinct individuals within each camera’s field of view.
+Algorithm 2 Occupancy Indicator Algorithm
+1:Input: data frame dfdf, historical_data
+2:for each camera_id in dfdf do
+3:  for every 5 seconds in record_time do
+4:   current_#_of_people = length(unique(global_IDs))
+5:  end for
+6:  if current_#_of_people $\leq$ percentile(historical_data, 25) then
+7:   occupancy = ”Low Occupancy”
+8:  else if current_#_of_people $\leq$ percentile(historical_data, 75) then
+9:   occupancy = ”Normal Occupancy”
+10:  else
+11:   occupancy = ”High Occupancy”
+12:  end if
+13:  update_historical_data(current_#_of_people)
+14:end for
+Occupancy levels are determined by comparing the current number of people against historical data percentiles.
+The algorithm classifies occupancy into three categories based on these thresholds: ’Low Occupancy’ when the count falls at or below the 25th percentile, ’Normal Occupancy’ when it is between the 25th and 75th percentiles, and ’High Occupancy’ when it exceeds the 75th percentile of historical data.
+Fig.
+5 compares occupancy indicators for the same camera during the same hour on two days.
+Graph 1 shows a weekend, with 2 people categorized as ”High Occupancy,” while Graph 2 shows the same number considered ”Normal” on a weekday.
+Following each occupancy determination, the update historical data function incorporates the current number of people into the historical dataset.
+This continuous updating mechanism ensures the algorithm’s thresholds remain adaptive and relevant, reflecting recent occupancy patterns.
+The entire process executes across all camera IDs in the data frame, providing comprehensive occupancy monitoring across the entire surveillance network.
+5.2.2 Statistical Anomaly
+Statistical Anomalies often indicate unusual traffic flow.
 By analyzing historical data, the system establishes a baseline, identifying deviations as anomalies.
-The system calculates such by comparing current detections with historical data, flagging events that exceed two standard deviations from the mean.
+The system calculates such anomalies by comparing current detections with historical data, flagging events that exceed two standard deviations from the mean.
 In our analysis, zero detections were excluded to avoid skewing the data.
-Bird’s Eye View
-provides an accurate spatial representation, ideal for crowd management and area planning.
+Algorithm 3 identifies statistically anomalous numbers of detected individuals for each camera relative to historical trends.
+Given that crowd densities vary across locations and times, these anomalies are calculated on an hourly basis.
+The approach involves constructing a normal distribution from historical hourly data, updated every 5 seconds as new detections occur.
+As outlined in Algorithm 3, the mean and standard deviation for each hour are calculated to characterize the data distribution.
+Current detections are then compared against this historical context.
+If the current number of detections exceeds two standard deviations from the mean, there is a less than 0.05 probability of such an occurrence, indicating with 95% confidence that it can be considered a statistical anomaly.
+Algorithm 3 Statistical Anomaly Detection
+1:Input: dfdf, start_timestart\_time, end_timeend\_time
+2: Filter dfdf for camera ID in the specified time range
+3: Initialize meanmean to 0 and stdstd to 1
+4: Initialize detected_objectsdetected\_objects list
+5:for each 5-second interval from start_timestart\_time to end_timeend\_time do
+6:  Count unique detected objects in the interval
+7:  if detected objects >>then
+8:   Update detected_objectsdetected\_objects, meanmean, and stdstd
+9:  end if
+10:end for
+We intentionally exclude detections of 0 in our analysis.
+The reason is that, within a 5-second window, the absence of detections (i.e., detecting no individuals) is more likely than detecting any individuals.
+Including 0 detections would skew the results toward zero, leading to lower thresholds for identifying anomalies and reducing the effectiveness of the statistical analysis.
+5.2.3 Bird’s Eye View
+Bird’s Eye View provides an accurate spatial representation, ideal for crowd management and area planning.
 The system normalizes object dimensions and uses a scale factor to accurately position objects within this view.
-Figure 5 compares the original and processed Bird’s Eye View from one camera, showing how this technique refines and clarifies spatial data.
-Heat map
-visually represent data distribution, useful for monitoring congestion and spatial usage.
+Algorithm 4 describes the transformation of camera-detected objects into Bird’s Eye View coordinates, specifically designed for the AXIS P3225-VE Mk II Network Camera [^1] https://www.axis.com/products/axis-p3225-ve-mk-ii/support#technical-specifications.
+This transformation creates a top-down perspective of the surveillance area.
+The transformation process begins with object dimension normalization, where width and height are divided by camera width and camera height, respectively.
+This step ensures dimensional independence from camera resolution.
+Subsequently, a scale factor is computed for each object based on its normalized height and the camera’s angular field of view parameters (min theta and max theta), which are obtained from the manufacturer’s specifications.
+This scaling mechanism preserves depth perception by appropriately adjusting object sizes based on their distance from the camera.
+The algorithm then calculates each object’s centroid by averaging its width and height dimensions, establishing a precise reference point for positional mapping.
+The final Bird’s Eye View coordinates (BirdsEyeX and BirdsEyeY) are derived by applying the computed scale factor to these centroid coordinates, resulting in an accurate top-down spatial representation.
+For temporal analysis, the algorithm implements a data filtering mechanism that processes records at 5-second intervals, tracking unique global IDs to determine object counts.
+This filtering ensures that analysis is performed only on relevant data within specified camera IDs and time frames.
+Algorithm 4 Bird’s Eye View Transformation
+1:Input: data frame dfdf, camera_width, camera_height, min_teta, max_teta
+2: Compute normalized values: $df[^{\prime}normalized\_W^{\prime}]$, $df[^{\prime}normalized\_H^{\prime}]$
+3:for each object in dfdf do
+4:  Compute scale_factorscale\_factor using normalized_Hnormalized\_H
+5:  Calculate centroid C_XC\_X and C_YC\_Y
+6:  Compute BirdsEye_XBirdsEye\_X and BirdsEye_YBirdsEye\_Y using scale_factorscale\_factor and normalized values
+7:end for
+8:for camera_idcamera\_id from 1 to 8 do
+9:  Define start_time and end_time
+10:  Filter dfdf based on camera_idcamera\_id and time range
+11:end for
+Fig.
+6 provides a comparative visualization between the original perspective and the processed Bird’s Eye View from one camera, captured simultaneously.
+In the initial view, each data point represents the XY coordinates of detected objects in the camera’s original field of view, showing approximately 12 distinct detections within a five-second interval.
+In the Bird’s Eye View, these XY coordinates are transformed, consolidating objects with the same global IDs into averaged positions.
+This results in nine distinct individuals, providing a top-down perspective that refines and simplifies the spatial arrangement.
+((a))
+((b))
+Figure 6: Comparing different views from one camera. Graph (a) generated directly from the database shows the original camera view. Graph (b) represents the averaged bird’s eye view coordinates.
+5.2.4 Heat map
+Heat maps visually represents data distribution, ann are useful for monitoring congestion and spatial usage.
 The system generates 2D heatmaps from Bird’s Eye View coordinates, applying Gaussian smoothing for clarity.
-Figure 6 presents heatmaps for different 24-hour intervals.
-Graph (a) and (c) show weekday pedestrian movement, while graph (b) and (d) represent the weekend congestions at the same areas.
-Comparing the density pattern during weekday and weekend, reflect differences in area usage.
-The detailed algorithms for all visualization tools are available in the supplementary materials for replication and further investigation.
-Community Engagement
-To gauge public perception of such solution, we conducted seven rounds of survey studies in six public places across Charlotte, NC, during July and August 2023.
+Fig.
+7 presents heatmaps for different 24-hour intervals.
+Graph (a) and (c) show weekday pedestrian movement, while Graph (b) and (d) represent the weekend congestion at the same areas.
+Comparing the density patterns during weekdays and weekends reflect differences in area usage.
+((a))
+((b))
+((c))
+((d))
+Figure 7: Heatmap representations from two cameras during weekdays and weekends.
+Graph (a) and (c) represent the heatmaps during weekdays, while Graph (b) and (d) show congestion in the same areas during weekends.
+Algorithm 5 is developed to generate heatmaps based on the data from cameras.
+We used the calculated Bird’s Eye View coordinates to generate heat maps.
+We processed data for each of the eight cameras, filtering records for every 24-hour window within the specified date range.
+Algorithm 5 Heatmap Generation
+1:Input: data frame dfdf, start_date, end_date
+2: Compute date_range from start_date to end_date
+3:for each date in date_range do
+4:  for camera_idcamera\_id from 1 to 8 do
+5:   Filter dfdf for current date and camera_idcamera\_id
+6:   if filtered data is not empty then
+7:    Determine max_xmax\_x and max_ymax\_y from filtered data
+8:    Initialize heatmap array of size (max_y+1,max_x+1)(max\_y+1,max\_x+1)
+9:    for each row in filtered data do
+10:     Increment heatmap value at position (Bird’s Eye $Y$, Bird’s Eye $X$)
+11:    end for
+12:   end if
+13:  end for
+14:end for
+The initial step involves defining a date range between a given start date and an end date.
+The algorithm then iterates over this date range and, for each date, processes data from each of the eight cameras.
+This systematic approach ensures comprehensive coverage, capturing the spatial dynamics of the monitored area over time.
+For each date and camera combination, the dataset, $df$, is filtered to include only records that fall within the current date and are associated with the current camera.
+Once filtered, the algorithm checks if there’s any data available for processing.
+If data is available for the given date and camera, the maximum Bird’s Eye View coordinates, max x and max y, are determined.
+These values define the dimensions of a 2D heatmap array, which is initialized with zeros.
+As the algorithm iterates over each row of the filtered data, the heatmap values are incremented based on the Bird’s Eye View positions of detected objects.
+Thus, by the end of this process, we obtain a series of heat maps, each capturing the spatial distribution of activity for a specific camera on a particular day, offering invaluable insights into the daily patterns of area usage and congestion.
+We applied Gaussian smoothing to the original heatmaps to enhance their clarity.
+6 Community Engagement
+To gauge public perception of our solution, we conducted seven rounds of survey studies in six public places across Charlotte, NC, during July and August 2023.
 A total of 410 respondents participated, providing insights into their concerns about safety and surveillance.
 The results indicated that respondents were most concerned about safety in parking lots (54%), public transit (44%), and entertainment venues (42%).
-Additionally, significant concerns were raised about potential biases and discrimination in current passive surveillance technologies, with these concerns being important or very important to respondents across all demographic groups, highlighting a widespread sensitivity to this issue.
+Additionally, significant concerns were raised about potential biases and discrimination in current passive surveillance technologies, being rated as important or very important to respondents across all demographic groups, highlighting a widespread sensitivity to this issue.
 [^2] This study has been previously published, but in adherence to the double-blind review policy of the AAAI conference, the reference will be provided in the camera-ready version upon acceptance.
@@
 These findings suggest a strong preference for AI-driven surveillance solutions over traditional passive technologies, especially in public places.
-System Evaluation and Results
-Figure 7: Latency and throughput trends concerning crowd densities during a week-long period for 8 camera nodes
-(a) Latency Trends
-(b) Throughput Trends
-Figure 8: Latency and Throughput trends with respect to crowd densities across different nodes number running in parallelism.
-Load Stress Evaluation
+7 System Evaluation and Results
+7.1 Load Stress Evaluation
 This section evaluated the system’s performance under increasing input loads by varying crowd densities from videos.
 Key metrics such as average latency and throughput were measured as we scaled the system using different numbers of pipeline nodes—one, four, eight, and twelve—to assess scalability and adaptability under increased workloads.
-To rigorously assess the system’s performance under varying crowd densities and parallel node operations, we prepared ten videos with density levels ranging from 0 to 9 from CHAD dataset (Pazho et al.
-2022).
-Each video lasted 150 seconds and operated at a frame rate of 30 frames per second.
-To ensure the statistical validity of our results, we averaged metrics over the central 100 batches in each run, excluding the initial 25 batches for system warm-up and the last 25 batches for cool-down effects.
-Figure 8 graphically presents the trends for throughput and latency.
-The X-axis, labeled ”Density” (count), represents the average number of humans detected per batch (comprising 30 frames) for each input node number in the experiment.
-With one and four camera nodes, our system maintained latency under 10 seconds, showcasing stability regardless of scene crowdness.
-However, with eight nodes, latency increased beyond 20 seconds at higher density levels, particularly from level 5 onwards.
-A notable decrease in latency was observed with 12 nodes, especially at density levels 8 and 9, correlating with processing 108 individuals (12 nodes x 9 density).
-Throughput consistently declined linearly with increasing density and node count, reaching a low of 4.56 FPS at the highest density level 9 with 12 nodes.
-The performance decrease is primarily due to the computational demands of pose estimator (Sun et al.
-2019) processing high-density keypoints and is further constrained by CPU computation limits.
-As node count and density increase, these factors collectively create a bottleneck, leading to the observed decline in throughput and latency.
-Real-World Endurance Evaluation
+Our performance evaluation methodology employed a systematic testing approach using the CHAD dataset [38], comprising ten videos with crowd density levels ranging from 0 to 9.
+Each test video was 150 seconds in duration with a 30-frames-per-second capture rate.
+To ensure robust statistical analysis and minimize the impact of system initialization and termination effects, we focused our analysis on the central 100 batches of each test run, excluding the first 25 batches (warm-up period) and the final 25 batches (cool-down period) from our calculations.
+((a))
+((b))
+Figure 8: Latency and Throughput trends with respect to crowd densities across different number of nodes running in parall.
+Graph (a) shows the latency trends, and Graph (b) shows the throughput trends.
+Figure 8 illustrates the relationship between system performance metrics and crowd density.
+The X-axis represents the average number of individuals detected per 30-frame batch across varying input node configurations.
+System performance exhibited distinct patterns across different node configurations.
+With one to four camera nodes, the system maintained stable latency below 10 seconds across all density levels.
+However, the eight-node configurations showed significant latency increases exceeding 20 seconds at density levels above 5.
+Interestingly, the twelve-node configuration demonstrated improved latency at high density levels (8 and 9), particularly when processing approximately 108 individuals.
+System throughput exhibited a linear decline correlating with increases in both density and node count, reaching its minimum of 4.56 FPS at density level 9 with 12 nodes.
+This performance degradation primarily stems from two factors: the computational intensity of the pose estimator [34] processing dense keypoint configurations, and CPU processing limitations.
+The combination of these factors creates a performance bottleneck that intensifies with increasing node count and crowd density.
+7.2 Real-World Endurance Evaluation
 This evaluation assessed the system’s long-term stability using all 16 cameras in the testbed.
-We conducted extended trials with eight, twelve, and sixteen nodes running continuously for 21 hours and a week-long test using eight nodes to ensure the system’s robustness.
+We conducted extended trials with eight, twelve, and sixteen nodes running continuously for 21 hours, and a week-long test using eight nodes to ensure the system’s robustness.
 This setup aimed to validate the system’s performance in realistic, extended operational conditions.
+  • 1.
     Eight-Node Setup: Covers key indoor areas, including entry points, hallways, and common spaces.
+  • 2.
     Twelve-Node Setup: Adds outdoor monitoring of parking lots.
+  • 3.
     Sixteen-Node Setup: Utilizes all cameras for comprehensive coverage.
-As shown in Figure 7, the system’s performance over a week, running with eight cameras, demonstrated consistent latency and throughput results about crowd density.
-Specifically, the latency ranged from 2.5-7.8s, and throughput varied from 22.3-29.8 FPS.
-These findings confirm that irrespective of crowd density, the latency increased.
-At the same time, the FPS decreased, aligning with expectations and underscoring that the length of the videos does not significantly impact the system’s performance.
-It is essential to note that the specific data spikes observed within this evaluation could stem from various factors, including network irregularities within the environment (as the input streams originate from IP cameras) or CPU memory usage nearing its limits.
-However, in light of the comprehensive experiment, these occasional data point spikes remain within acceptable bounds.
-The system consistently maintained optimal performance levels, showcasing its robustness and capacity to effectively manage and adapt to various operational scenarios.
-For configurations with 8, 12, and 16 camera nodes, we observed similar trends over the 21-hour-long experiment with varying performance metrics: average latencies ranged from 2.6-4.8s, 5.3-6.5s, and 6.7-10.5s, respectively, while corresponding throughputs ranged from 28.5-26.5 FPS, 20.5-18 FPS, and 16.5-14.5 FPS.
-Detailed results with visualizations for 8, 12, and 16 camera nodes are provided in the supplementary materials for replication and further investigation.
-(a) Object Anomaly
-(b) Behavior Anomaly
-Figure 9: Represent latency rug plot with different anomaly scenarios. 50 data points for each camera number
-Physical-Cyber-Physical Evaluation (Anomaly detection)
+Figures 9, 10, and 11 present comprehensive analyses of system latency and throughput patterns across varying crowd densities during a 21-hour operational period.
+The evaluation, conducted from 2:00 p.m.
+to 11:00 a.m.
+the following day, was tested with three camera node configurations.
+Our observations revealed a notable decrease in system latency when the server node’s database approached 50,000 queries.
+To maintain optimal performance and ensure data privacy, we implemented an automatic database reset mechanism that activates either upon reaching 50,000 queries or after 24 hours of continuous operation.
+Figure 9: Latency and throughput trends concerning crowd densities during a 21-hour-length period for 8 camera nodes
+Figure 10: Latency and throughput trends concerning crowd densities during a 21-hour-length period for 12 camera nodes
+Figure 11: Latency and throughput trends concerning crowd densities during a 21-hour-length period for 16 camera nodes
+The data presented in Figures 9, 10, and 11 represent values averaged over 60-second intervals.
+The ’Density’ metric reflects the aggregate number of human detections per camera node during these one-minute sampling windows.
+The relatively low human traffic observed, particularly from indoor camera nodes, can be attributed to the data collection period coinciding with the summer break period.
+Figure 12: Latency and throughput trends concerning crowd densities during a week-long period for 8 camera nodes
+Figure 12 illustrates the system’s week-long performance using an eight-camera configuration.
+The results demonstrate consistent correlations between crowd density and performance metrics, with latency ranging from 2.5-7.8 seconds and throughput varying between 22.3-29.8 FPS.
+These findings confirm our expected inverse relationship between crowd density and system performance, while demonstrating that video duration has minimal impact on overall system performance.
+Within the dataset, occasional performance spikes were observed, potentially attributable to environmental network fluctuations from IP camera streams or periods of peak CPU memory utilization.
+However, these variations remained within acceptable operational parameters, demonstrating the system’s resilience under diverse conditions.
+The extended 21-hour evaluation across different camera configurations revealed consistent performance patterns:
+  • 1.
+    8 nodes: latency of 2.6-4.8s with throughput of 28.5-26.5 FPS
+  • 2.
+    12 nodes: latency of 5.3-6.5s with throughput of 20.5-18 FPS
+  • 3.
+    16 nodes: latency of 6.7-10.5s with throughput of 16.5-14.5 FPS
+The system demonstrated consistent performance under diverse operational conditions, validating its robustness and adaptability.
+7.3 Physical-Cyber-Physical Evaluation (Anomaly detection)
 This section evaluates the system’s Physical-Cyber-Physical (PCP) latency, representing the end-to-end time from when an anomaly appears to when the end user receives a notification.
-We assessed both object and behavioral anomaly detection (Markovitz et al.
-2020b) using four, eight, twelve, and sixteen AI pipeline nodes reading input frames from the same camera node, conducting twenty experiments per anomaly type and collecting 400 data samples.
-The latency measures the time from detecting anomalies, such as high-priority objects or dangerous behaviors like fighting, to alerting the end user.
-A three-second transmission delay from cameras to AI Nodes was observed, slightly higher than other IP systems, due to our unique network setup.
+To showcase the versatility and robustness of our AI-driven surveillance system, we established Testbed B with 12 CCTV cameras—3 outdoor and 9 indoor—at a busy market shop, featuring a 32-core CPU, 512 GB RAM, and two 48 GB VRAM GPUs.
+This setup complements Testbed A, providing diverse environmental conditions and hardware configurations for testing anomaly detection.
+For a comparative baseline, we deployed the same pipeline on AWS cloud, using an EC2 p2.xlarge instance with a single GPU to process video streams from IoT IP cameras over Real-Time Streaming Protocol.
+Metadata is stored in Amazon RDS, queried via Athena, and notifications are sent through SNS, mirroring the edge architecture for a direct performance comparison.
+Fig.
+13 illustrates example camera views from two testbeds without any human presence.
+By comparing the edge-based setups with the cloud deployment, we specifically examined end-to-end latency differences.
+In the cloud setup, the EC2 instance consolidates local and global processing nodes, with Amazon services enabling real-time analytics and alerts.
+This configuration highlights how offloading inference to the cloud impacts response times, with results from the p2.xlarge instance serving as a reference point.
+While alternative AWS configurations could yield different latencies and costs, this setup provides valuable insights into the trade-offs of cloud processing, aiding in the optimization of surveillance system design based on operational needs.
+Figure 13: Example footages of the cameras in different locations
+We assessed both object and behavioral anomaly detection [29] using four, eight, twelve, and sixteen AI pipeline nodes reading input frames from the same camera node, conducting fifty experiments per anomaly type and collecting 700 data samples.
+PCP latency measures the time from detecting anomalies, such as high-priority objects or dangerous behaviors like fighting, to alerting the end user.
+As mentioned in Sec 3, the object detector can detect object-of-interest classes as object anomaly and behavior anomaly detector is trained on normal movement sequences and evaluates incoming sequences to see whether the movement is normal or not.
+For evaluation, we used predefined test cases: prohibited items were placed in predetermined locations, and staged physical confrontations were enacted by trained, authorized participants within the detection range of the camera under examination.
+These controlled scenarios ensured consistent triggering of the system’s anomaly detection mechanisms, enabling accurate measurement of end-to-end alert latency.
+Various strategies were implemented to reflect the system’s versatility across the two testbeds.
+For testbed A, although the number of camera nodes was increased, multiple streams were duplicated from the same camera, ensuring that crowd density remained controlled.
+By contrast, testbed B, employed different camera streams onsite when increasing the node count, resulting in an uncontrolled number of individuals entering the field of view, thereby simulating real-world conditions.
+About three-second transmission delay from cameras to AI nodes was observed for testbed A, slightly higher than other IP systems, due to its unique network setup.
 This delay is included in the overall PCP measurements, though manual recording of latency data may introduce minor discrepancies.
+Table 4 provides more detailed insight about P-C-P evaluation results.
+This evaluation demonstrates the system’s capability to enhance safety and operational efficiency in real-world scenarios.
+The ability to notify end users within approximately 20 seconds enables transforming the existing CCTV systems from passive video cameras to proactive systems that can notify the stakeholders about the events that are happening.
 The results indicate that increasing the number of cameras significantly impacts PCP latency.
-For example, object anomaly latency increased from 4.7 seconds with four cameras to 19.68 seconds with sixteen, while behavioral anomaly latency rose from 9.34 seconds to 26.76 seconds.
-This increase is attributed to system traffic, multitasking overhead, and the processing capacity of the system.
-Figure 9 displays the distribution of 400 data points from our PCP experiments.
-Graph (a) shows the distribution for object anomaly detection, with less variation in the four-camera setup compared to the sixteen-camera setup.
-Graph (b) illustrates the distribution for behavioral anomaly detection, where the data is more scattered with an increased number of cameras.
-The standard deviation for object detection latency increased from 0.65 to 1.12 seconds as the camera count increased, indicating a slight broadening of latency due to system complexity and resource contention.
-This evaluation demonstrates the system’s capability to enhance safety and operational efficiency in real-world scenarios.
-The ability to notify end users within approximately 20 seconds is crucial for managing anomalies in public spaces, aiding in the prevention of potential dangers and improving resource allocation.
-Overall, these real-world evaluations allows research how these key parameters, such as the number of cameras and crowd density, could affect system performance, such as latency and throughput in different public safety scenarios in case of emergency, is crucial in decision making and taking the proper actions.
-For example, retail and transport hubs could analyze foot traffic and crowd density, offering insights into peak hours and emergency response planning (Hosseini et al.
-2024).
-Conclusion
+For example, object anomaly latency increased from 4.7 seconds with four cameras to 19.68 seconds with sixteen, while behavioral anomaly latency rose from 9.34 seconds to 26.76 seconds in testbed A.
+This increase is attributed to higher system traffic, multitasking overhead, and the overall processing capacity.
+Tests at Testbed B confirmed these trends, illustrating how a more dynamic environment and increased workload can further affect end-to-end alert times.
+In contrast, the cloud-based system offers significant advantages in terms of performance stability.
+Unlike the edge system, which is constrained by its local processing capabilities and experiences substantial latency increases as more cameras are added, the cloud infrastructure can dynamically allocate additional processing instances to handle the increased workload from more cameras.
+However, the result of 21.38 seconds for object anomaly detection and 26.5 seconds for behavior anomaly detection is not competitive with the edge based system in the experiment.
+Table 4: Statistical PCP latency data for object and behavior anomaly for different number of cameras in different environment
+Location A (Controlled) Location B (Open) Full Cloud
+Nodes Mean (s) Min (s) Max (s)
+Standard
+deviation
+Mean (s) Min (s) Max (s)
+Standard
+deviation
+Mean(s)
+Object Anomaly 4 4.7 3.06 6.4 0.55 3.372 2.29 5.01 0.64 21.38
+8 10.99 9.18 12.55 0.85 8.924 7.46 10.61 0.71
+12 15.76 14.05 18.31 1.05 13.037 10.28 16.93 1.27
+16 19.68 17.67 21.6 1.13 - - - -
+Behavior Anomaly 4 9.34 7.81 10.41 0.65 6.09 4.78 7.37 0.6 26.5
+8 14.53 13.23 16.46 0.71 11.62 10.75 13.3 0.62
+12 18.45 16.32 20.35 0.97 17.68 15.42 20.55 1.06
+16 26.76 24.23 29.54 1.52 - - - -
+Overall, these real-world evaluations allow research into how key parameters—such as the number of cameras and crowd density—affect system performance metrics like latency and throughput.
+Understanding these dynamics in different public safety scenarios, especially in cases of emergency, is crucial for informed decision-making and timely action.
+For instance, in retail environments and transportation hubs, the system can provide valuable insights into foot traffic patterns and crowd density fluctuations, facilitating both operational optimization and emergency response.
+[39]
+8 Conclusion
 This paper presented a comprehensive evaluation of an AI-based real-time video solution, integrated with existing CCTV infrastructures, to enhance situational awareness through object and behavioral anomaly detection.
@@
 Object and behavioral anomaly detection latencies ranged from 4.7 seconds with four cameras to 26.76 seconds with sixteen, reflecting the system’s scalability and robustness.
-Through extensive testing, the system demonstrated its capability to deliver timely notifications to end users, crucial for managing anomalies in public spaces.
+Through extensive testing, the system demonstrated its capability to deliver timely notifications to end users, which is crucial for managing anomalies in public spaces.
 This work underscores the importance of real-world evaluations in optimizing AI-driven video solutions, ensuring they meet the demands of dynamic environments and contribute to enhanced public safety and operational efficiency.